OpenStudy (shamil98):

You have a coupon for your favorite clothing store for $20 off any purchase of more than $50. The store is also running a 25%-off sale on its entire inventory. Let x be the original price, f(x)be the price with the $20 coupon applied, and g(x) be the price with the 25% discount applied. Find f(x) and g(x), how would i write them? c. What would the expression (f o g) (x) represent? d. c. What would the expression (g o f) (x) represent?

OpenStudy (anonymous):

Alright its oing to take me about 10 mins whats ur options in answers?

OpenStudy (shamil98):

There aren't any. lol. Free response.

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

working on it now

OpenStudy (anonymous):

x*g(x)-f(x)

OpenStudy (anonymous):

the orginal price multiplied by 25% off minus the 20 dollor coupon

OpenStudy (anonymous):

Wouldn't it be \[f(x)=x-20;x \ge50\] and \[g(x)=f(x)*0.75\]Sorry if I made a mistake, its late at night.